# Hybrid collective excitations in topological superconductor/ferromagnetic insulator heterostructures

### v3

@@ Abstract
We develop a linear response theory for the dynamical proximity effect in topological superconductor/ferromagnetic insulator (TS/FI) hybrid structures. Our approach combines the nonequilibrium quasiclassical Keldysh–Usadel equations for the electronic Green’s functions in the TS with the Landau–Lifshitz–Gilbert equation governing the magnetization dynamics in the FI. Within this framework, we study the proximity-induced coupling between magnons and superconducting collective excitations. We find that the spin–momentum locking intrinsic to the surface state of the TS leads to a hybridization between the superconducting Nambu–Goldstone (phase) collective mode and magnons, resulting in the emergence of composite magnon–Nambu–Goldstone excitations. The dependence of the coupling strength on relevant physical parameters is analyzed both analytically and numerically. In contrast, we show that the Higgs (amplitude) mode does not couple to magnons at linear order and therefore does not participate in the formation of hybrid collective excitations.
@@ I Introduction
On the other hand, when the system becomes time-dependent, the dynamical properties of the magnetic and superconducting subsystems can be significantly modified in S/F hybrid structures [2]. For instance, due to partial singlet–triplet conversion, composite quasiparticles consisting of magnons and clouds of triplet Cooper pairs can emerge at the interface between a superconductor and a ferromagnetic (FI) or antiferromagnetic (AFI) insulator [14, 15]. In addition to the partial conversion of singlet Cooper pairs into triplet states via proximity to a magnetic system, the superconducting condensate possesses another important property: the ability to sustain collective modes.
@@
Typically, the phase mode is excluded from the analysis in three-dimensional (3D) superconductors, since it is lifted up to the plasma frequency in the presence of Coulomb interaction and becomes indistinguishable from plasma oscillations[21]. However, the situation differs fundamentally in two-dimensional (2D) systems, where plasmon dispersion becomes gapless[47]. This implies that when such a system enters the superconducting state, the associated phase mode should also remain gapless[48]. Importantly, in 3D case in the vicinity of the superconducting critical temperature, charge neutrality can be maintained due to a large population of quasiparticles. This condition allows for the emergence of a gapless collective excitation, known as the Carlson-Goldman (CG) mode [49, 50, 51, 52, 31, 53]. Moreover, the interplay between plasmon and CG mode in thin superconductors was also examined theoretically[54]. Thus, considering the rapid advancements of superconducting 2D materials and topological surface physics [55, 56], exploring the phase mode in 2D systems becomes essential for further understanding of fundamental physical mechanisms as well as possible applications in the field of low dissipation electronics. Despite the achieved progress in the studies of the dynamical proximity effects, the coupling between magnons and superconducting phase Nambu-Goldstone (NG) modes have not yet been reported in the literature.
@@
In this work we focus on the examination of the mutual influence of the superconducting collective modes and magnon excitations in topological superconductor/ferromagnetic insulator (TS/FI) hybrid structures. The conductive surface state of the TS represents a basic system sustaining the strongest SOC in the form of the full spin-momentum locking [57, 58, 59, 60]. Due to the 2D nature of this superconducting state, the NG mode is not lifted up to the plasma frequency even if the Coulomb interaction effects are taken into account and remains a soft mode [48, 61, 62, 31]. The interaction of magnons with surface plasmons of magnetic insulators and nonsuperconducting topological insulators was already discussed[63]. Here, we develop a linear response theory of collective excitations in 2D superconducting systems having the property of the full spin-momentum locking contacted with a thin-film ferromagnetic insulator. The theory is developed in the framework of the non-equilibrium Keldysh quasiclassical approach. We show that the spin-momentum locking leads to hybridization of the NG phase mode with the magnon and the appearance of composite excitations consisting of a magnon in the FI accompanied by oscillations of the phase of the superconducting order parameter in TS. The dependence of the coupling strength on the relevant physical parameters is studied analytically and numerically. At the same time, it is demonstrated that the Higgs mode is not coupled to magnons in the linear order and does not form hybrid collective excitations with it.
@@
The paper is organized as follows. In Sec. II we formulate the studied model. In Sec. III all stages of theory construction are described. In Sec. IV all results for the spectra of uncoupled and coupled superconducting modes and magnons obtained on the basis of the developed theory are presented. Our conclusions are summarized in Sec. V. In Appendix A some details of the Green’s functions calculations are given and in Appendix B some symmetry relations between the different components of the Green’s function are presented, which are used to prove the absence of interaction between the magnon and the Higgs mode.
@@ II System and model
Figure 1: Schematic sketch of the TS/FI hybrid heterostructure. Plots (a), (b) and (e) correspond to the side view of the system, FI and TS top views, respectively. The superconducting and magnetic subsystems interact via interface exchange coupling constant $J_{ex}$, which is illustrated in (a). Panel (d) shows the normal state dispersion relation of the 2D conductive surface state of the TI with spin-momentum locking. The basic modes of the non-interacting system (i.e. $J_{ex}=0$) are shown in plot (c).
We consider a topological superconductor/ferromagnetic insulator (TS/FI) heterostructure, see Fig. 1. On the surface of the TS, a 2D superconducting state with full spin-momentum locking occurs. The corresponding normal state electron dispersion is shown in Fig. 1(d). The normal state Fermi surface of the conductive surface state of the TS is represented by the only helical band [64, 65] similar to the conductive surface state of topological insulator [57, 58, 59, 60]. The electron spin at the Fermi surface always makes the right angle with its momentum with a definite helicity.
@@
$H$ $=\int d^{2}r\Biggl{\{}\Psi^{\dagger}(\bm{r})\bigl{[}-i\hbar v_{f}(\bm{\nabla}_{\bm{r}}\times\hat{z})\bm{\sigma}$
$-\mu+e\phi(\bm{r})+V_{imp}(\bm{r})-\bm{h}\bm{\sigma}\bigr{]}\Psi(\bm{r})$
$+\Delta(\bm{r})\Psi^{\dagger}_{\uparrow}(\bm{r})\Psi^{\dagger}_{\downarrow}(\bm{r})+\Delta^{*}(\bm{r})\Psi_{\downarrow}(\bm{r})\Psi_{\uparrow}(\bm{r})\Biggr{\}},$~~~~ (1)
@@
where $\Psi^{\dagger}(\bm{r})=(\Psi^{\dagger}_{\uparrow}(\bm{r}),\Psi^{\dagger}_{\downarrow}(\bm{r}))$ is the electron creation operator, $\hat{z}$ is the unit vector normal to the surface of TS, $v_{f}$ is the electron Fermi velocity, $\mu$ is the chemical potential, $\phi(\bm{r})$ is the scalar electric potential and $\Delta(\bm{r})$ is the superconducting order parameter (OP), which is assumed to be of $s$-wave singlet type and should be calculated self-consistently as $\Delta(\bm{r})=\lambda\langle\Psi_{\downarrow}\Psi_{\uparrow}\rangle$. $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is a vector of Pauli matrices in spin space. The term $V_{imp}(\bm{r})$ includes the nonmagnetic impurity scattering potential $V_{imp}(\bm{r})=\sum\limits_{\bm{r}_{i}}V_{i}\delta(\bm{r}-\bm{r}_{i})$, which is of a Gaussian form $\langle V_{imp}(\bm{r})V_{imp}(\bm{r}^{\prime})\rangle=(1/\pi\nu\tau)\delta(\bm{r}-\bm{r}^{\prime})$ with $\nu=\mu/(2\pi v_{f}^{2})$.
@@
The magnetization in the FI takes the form $\bm{m}(\bm{r},t)=\bm{m}_{0}+\bm{\delta m}(\bm{r},t)$, where $\bm{m}_{0}=\bm{\hat{}}x$ is the equilibrium magnetization and $\bm{\delta m}(\bm{r},t)={\rm Re}[\delta\bm{m}e^{i(\bm{k\cdot r}+\omega t)}]\exp{(-\kappa t)}$ accounts for the spin wave, where $\kappa$ characterizes the decay of the spin wave. Then the induced in the TS exchange field takes the form $\bm{h}=\bm{h}_{0}+\delta\bm{h}$, where $\bm{h}_{0}=-J_{ex}M_{s}\bm{m}_{0}/(2\gamma d_{s})=h_{0}\hat{x}$ is the equilibrium exchange field. We consider the limit $\mu\gg(h,\Delta)$, when the Fermi level of the conductive surface state of TS is far from the Dirac point. In this limit the out-of-plane $z$-component of the magnon excitation produces negligible $z$-component of the effective exchange field, which is $\sim(h_{0}/\mu)m_{z}$ [72, 73], and we disregard it. Then the magnon-induced effective exchange field $\delta\bm{h}$ takes the form
$\delta\bm{h}=\delta h\cos\left(\omega t+\bm{k}\cdot\bm{r}\right)e^{-\kappa t}\hat{y}$=
$(\delta\bm{h}_{\omega,\bm{k}}e^{i\left(\omega t+\bm{k}\cdot\bm{r}\right)}+\delta\bm{h}_{-\omega,-\bm{k}}e^{-i\left(\omega t+\bm{k}\cdot\bm{r}\right)})e^{-\kappa t},$ (2)
@@
where $\delta\bm{h}_{\omega,\bm{k}}=\delta\bm{h}_{-\omega,-\bm{k}}=(\delta h/2)\hat{y}$. If one assumes a spatially homogeneous order parameter in the equilibrium state of the system $\Delta(\bm{r})=\Delta$, then in this state a spontaneous electric current occurs [74]. It was shown that such an equilibrium state does not have a minimal energy. The ground state of the system in the presence of the equilibrium exchange field $\bm{h}_{x}$ is the so-called helical state [75, 76, 77, 78, 79, 80, 81] and is described by the phase-inhomogeneous order parameter $\Delta(\bm{r})=\Delta\exp[i\bm{q}\bm{r}]$, where $\Delta$ is the absolute value of the OP at a given temperature $T$, and $\bm{q}=-2h_{0}\hat{e}_{y}/\hbar v_{f}$ is determined from the condition that the total current in the ground state is zero.
III Theoretical approach to the calculation of hybrid collective excitations

### v4

@@ Abstract
We develop a linear response theory for the dynamical proximity effect in topological superconductor/ferromagnetic insulator (TS/FI) hybrids. Our approach integrates the nonequilibrium quasiclassical Keldysh–Usadel formalism for the TS with the Landau–Lifshitz–Gilbert equation for the FI’s magnetization dynamics. This framework reveals a proximity-induced coupling between magnons and superconducting collective modes. Crucially, we find that spin–momentum locking in the TS surface state drives a hybridization between magnons and the superconducting Nambu–Goldstone (phase) mode, giving rise to composite magnon–Nambu-Goldstone excitations. We analyze the coupling strength’s dependence on key parameters both analytically and numerically. In contrast, we demonstrate that the Higgs (amplitude) mode does not couple to magnons at linear order and is thus excluded from the hybrid excitation spectrum. The hybridization between magnons and the superconducting phase mode provides a mechanism for the interconversion of spin signals and the spinless signals carried by collective superconducting excitations, thereby giving new impetus to the development of superconducting spintronics.
@@ I Introduction
On the other hand, when the system becomes time-dependent, the dynamical properties of the magnetic and superconducting subsystems can be significantly modified in S/F hybrid structures [36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54]. For instance, due to partial singlet–triplet conversion, composite quasiparticles consisting of magnons and clouds of triplet Cooper pairs can emerge at the interface between a superconductor and a ferromagnetic (FI) or antiferromagnetic (AFI) insulator [55, 56]. In addition to the partial conversion of singlet Cooper pairs into triplet states via proximity to a magnetic system, the superconducting condensate possesses another important property: the ability to sustain collective modes.
@@
Typically, the phase mode is excluded from analysis in three-dimensional (3D) superconductors. This is because, in the presence of long-range Coulomb interactions, the phase mode hybridizes with the gapped 3D plasmon. This hybridization lifts the phase mode’s energy up to the plasma frequency, making it indistinguishable from plasma oscillations [62]. An important exception occurs near the superconducting critical temperature, where a large population of thermally-excited quasiparticles can screen the charge. This allows a gapless phase mode, known as the Carlson-Goldman (CG) mode, to emerge [88, 89, 90, 91, 72, 92]. The interplay between plasmons and the CG mode in thin films has been a subject of theoretical study [93].
@@
Crucially, the situation is fundamentally different in two-dimensional (2D) systems. In 2D, the plasmon itself becomes a gapless mode, with a dispersion that vanishes as momentum goes to zero [94]. Therefore, when this gapless 2D plasmon hybridizes with the superconducting phase mode, it cannot impart a gap. Consequently, the phase mode in a 2D superconductor remains gapless at all temperatures [95, 96, 97, 72, 98]. Thus, considering the rapid advancements of superconducting 2D materials and topological surface physics [99, 100], exploring the phase mode in 2D systems becomes essential for further understanding of fundamental physical mechanisms as well as possible applications in the field of low dissipation electronics. Despite the achieved progress in the studies of the dynamical proximity effects, the coupling between magnons and superconducting phase Nambu-Goldstone (NG) modes have not yet been reported in the literature.
@@
In this work we focus on the examination of the mutual influence of the superconducting collective modes and magnon excitations in topological superconductor/ferromagnetic insulator (TS/FI) hybrid structures. The conductive surface state of the TS represents a basic system sustaining the strongest SOC in the form of the full spin-momentum locking [101, 102, 103, 104]. The interaction of magnons with surface plasmons of magnetic insulators and nonsuperconducting topological insulators was already discussed [105, 106]. Furthermore, the interaction of the superconducting plasmon in a one-dimensional resonator and magnons in an antiferromagnetic insulator was also discussed theoretically [107]. Here, we develop a linear response theory of collective excitations in 2D superconducting systems having the property of the full spin-momentum locking contacted with a thin-film ferromagnetic insulator. The theory is developed in the framework of the non-equilibrium Keldysh quasiclassical approach. We show that the spin-momentum locking leads to hybridization of the NG phase mode with the magnon and the appearance of composite excitations consisting of a magnon in the FI accompanied by oscillations of the phase of the superconducting order parameter in TS. The dependence of the coupling strength on the relevant physical parameters is studied analytically and numerically. At the same time, it is demonstrated that the Higgs mode is not coupled to magnons in the linear order and does not form hybrid collective excitations with it.
@@
The paper is organized as follows. Section II formulates the studied model. Section III presents a qualitative picture of the effect and details the construction of the theoretical framework. The resulting spectra of both uncoupled and coupled superconducting modes and magnons are presented in Sec. IV, and our conclusions are summarized in Sec. V. Further supporting details are provided in the appendices: Appendix A covers technical aspects of the Green’s function calculations, and Appendix B presents symmetry relations used to prove the absence of magnon-Higgs mode interaction.
@@ II System and model
Figure 1: Schematic of the TS/FI heterostructure. (a) Side view of the system, illustrating the interface exchange coupling $J_{ex}$ between the superconducting and magnetic subsystems. Normal-state dispersion relation of the 2D conductive surface state of the TI, showing spin-momentum locking is depicted at the top right corner. The equilibrium magnetization of the FI $\bm{m}_{0}=\hat{x}$ and a magnon excitation $\delta\bm{m}$ of the FI induce in the TS an equilibrium effective exchange field $\bm{h}_{x}$ and a magnon-induced dynamic contribution $\delta\bm{h}_{y}$, respectively. The dynamic exchange field $\delta\bm{h}_{y}$ triggers a time-dependent response of the superconducting order parameter in the TS, resulting in a hybridized magnon–Nambu–Goldstone (NG) excitation. (b) Top view of the FI layer. (c) Top view of the TS layer. (d) The basic modes of the non-interacting system (i.e. $J_{ex}=0$). (e) Schematic plot of the magnon -NG mode hybridization in the system and (f) anticrossing strength as a function of the temperature.
We consider a topological superconductor/ferromagnetic insulator (TS/FI) heterostructure, see Fig. 1(a)-(c). On the surface of the TS, a 2D superconducting state with full spin-momentum locking occurs. The corresponding normal state electron dispersion is shown in Fig. 1(a). The normal state Fermi surface of the conductive surface state of the TS is represented by the only helical band [108, 109] similar to the conductive surface state of topological insulator [101, 102, 103, 104]. The electron spin at the Fermi surface always makes the right angle with its momentum with a definite helicity.
@@
$H$ $=\int d^{2}r\Biggl\{\Psi^{\dagger}(\bm{r},t)\bigl[-i\hbar v_{f}(\bm{\nabla}_{\bm{r}}\times\hat{z})\bm{\sigma}$
$-\mu+e\phi(\bm{r})+V_{imp}(\bm{r})-\bm{h}\bm{\sigma}\bigr]\Psi(\bm{r},t)$
$+\Delta(\bm{r},t)\Psi^{\dagger}_{\uparrow}(\bm{r},t)\Psi^{\dagger}_{\downarrow}(\bm{r},t)+\Delta^{*}(\bm{r},t)\Psi_{\downarrow}(\bm{r},t)\Psi_{\uparrow}(\bm{r},t)\Biggr\},$~~~~ (1)
@@
where $\Psi^{\dagger}(\bm{r},t)=(\Psi^{\dagger}_{\uparrow}(\bm{r},t),\Psi^{\dagger}_{\downarrow}(\bm{r},t))$ is the electron creation operator, $\hat{z}$ is the unit vector normal to the surface of TS, $v_{f}$ is the electron Fermi velocity, $\mu$ is the chemical potential, $\phi(\bm{r})$ is the scalar electric potential and $\Delta(\bm{r},t)$ is the superconducting order parameter (OP), which is assumed to be of $s$-wave singlet type and should be calculated self-consistently as $\Delta(\bm{r},t)=\lambda\langle\Psi_{\downarrow}\Psi_{\uparrow}\rangle$. $\bm{\sigma}=(\sigma_{x},\sigma_{y},\sigma_{z})$ is a vector of Pauli matrices in spin space. The term $V_{imp}(\bm{r})$ includes the nonmagnetic impurity scattering potential $V_{imp}(\bm{r})=\sum\limits_{\bm{r}_{i}}V_{i}\delta(\bm{r}-\bm{r}_{i})$, which is of a Gaussian form $\langle V_{imp}(\bm{r})V_{imp}(\bm{r}^{\prime})\rangle=(1/\pi\nu\tau)\delta(\bm{r}-\bm{r}^{\prime})$ with $\nu=\mu/(2\pi v_{f}^{2})$.
Our goal is to consider the dynamical coupling between the magnons and the superconductive collective modes. For this reason the magnetization in the FI is assumed to be time-dependent and takes the form $\bm{m}(\bm{r},t)=\bm{m}_{0}+\bm{\delta m}(\bm{r},t)$, where $\bm{m}_{0}=\bm{\hat{}}x$ is the equilibrium magnetization and $\bm{\delta m}(\bm{r},t)={\rm Re}[\delta\bm{m}e^{i(\bm{k\cdot r}+\omega_{r}t)}]\exp{(-\kappa t)}$ accounts for the spin wave, where $\omega_{r}$ is the magnon frequency and $\kappa$ characterizes the decay of the spin wave, which originates from the intrinsic damping of the magnet and also from the coupling to the superconducting excitations. $\bm{m}(\bm{r},t)$ obeys the Landau-Lifshits-Gilbert equation (LLG), see Sec. III.3.
@@
The induced in the TS exchange field takes the form $\bm{h}=\bm{h}_{0}+\delta\bm{h}$, where $\bm{h}_{0}=-J_{ex}M_{s}\bm{m}_{0}/(2\gamma d_{s})=h_{0}\hat{x}$ is the equilibrium exchange field. We consider the limit $\mu\gg(h,\Delta)$, when the Fermi level of the conductive surface state of TS is far from the Dirac point. In this limit the out-of-plane $z$-component of the magnon excitation produces negligible $z$-component of the effective exchange field, which is $\sim(h_{0}/\mu)m_{z}$ [115, 116], and we disregard it. Then the magnon-induced effective exchange field $\delta\bm{h}$ takes the form
$\delta\bm{h}=\delta h\cos\left(\omega_{r}t+\bm{k}\cdot\bm{r}\right)e^{-\kappa t}\hat{y}$=
$(\delta\bm{h}_{\omega_{r},\bm{k}}e^{i\left(\omega_{r}t+\bm{k}\cdot\bm{r}\right)}+\delta\bm{h}_{-\omega_{r},-\bm{k}}e^{-i\left(\omega_{r}t+\bm{k}\cdot\bm{r}\right)})e^{-\kappa t},$ (2)
@@
where $\delta\bm{h}_{\omega_{r},\bm{k}}=\delta\bm{h}_{-\omega_{r},-\bm{k}}=(\delta h/2)\hat{y}$. If one assumes a spatially homogeneous order parameter in the equilibrium state of the system $\Delta(\bm{r})=\Delta$, then in this state a spontaneous electric current occurs [117]. It was shown that such an equilibrium state does not have a minimal energy. The ground state of the system in the presence of the equilibrium exchange field $\bm{h}_{x}$ is the so-called helical state [118, 119, 120, 121, 122, 123, 124] and is described by the phase-inhomogeneous order parameter $\Delta(\bm{r})=\Delta\exp[i\bm{q}\bm{r}]$, where $\Delta$ is the absolute value of the OP at a given temperature $T$, and $\bm{q}=-2h_{0}\hat{e}_{y}/\hbar v_{f}$ is determined from the condition that the total current in the ground state is zero. The superconducting collective modes are described by the time dependent correction to the OP, that is in general $\Delta(\bm{r},t)=\Delta\exp[i\bm{q}\bm{r}]+\delta\Delta(\bm{r},t)$.
III Qualitative picture and theoretical formalism
This section outlines the qualitative behavior of collective excitations and presents the theoretical framework for calculating their spectrum in the TS/FI hybrid system, based on the non-equilibrium quasiclassical Keldysh formalism.
The spectrum is determined by coupled fluctuations of the magnetization in the FI, the complex-valued order parameter in the TS, and the scalar potential, the latter being necessary to account for charge oscillations. The observables corresponding to these fluctuating fields are derived from the Keldysh Green’s functions in the linear response approximation, as detailed in Sec. III.1. The dynamics are governed by three key equations: the self-consistency equation for superconducting gap oscillations ( Sec. III.1), the Poisson equation for the self-consistent scalar potential under Coulomb interaction (Sec. III.2), and the LLG equation for magnon dynamics in the FI (Sec. III.3). These equations must be solved simultaneously to obtain the full spectrum as well as damping (decay rate) of the collective modes (Sec.III.4), as summarized in Table 1.
The spin-momentum locking provides a linear response of the OP to $\delta\bm{h}$:
$\delta\Delta(\bm{r},t)=(\delta{\Delta}_{\omega_{r},k}e^{i\left(\omega_{r}t+\bm{k}\cdot\bm{r}\right)}+$
$\delta{\Delta}_{-\omega_{r},-k}e^{-i\left(\omega_{r}t+\bm{k}\cdot\bm{r}\right)})e^{-\kappa t}.$~~~~ (3)
It is convenient to decompose the dynamical correction to the OP into contributions from the amplitude (Higgs) mode and the phase (NG) mode, defined respectively as $\delta\Delta_{a}=[\delta\Delta_{\omega_{r},\bm{k}}+\delta\Delta_{-\omega_{r},-\bm{k}}^{*}]/2$ and $\delta\Delta_{p}=[\delta\Delta_{\omega_{r},\bm{k}}-\delta\Delta_{-\omega_{r},-\bm{k}}^{*}]/2i$. The corresponding linear response corrections can then be written in the form:
$\delta\Delta^{a}=F_{\omega_{r},k}^{a}(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\hat{z}$,
$\delta\Delta^{p}=F_{\omega_{r},k}^{p}(\hat{\bm{k}}\times\delta\bm{h}_{\omega_{r},\bm{k}})\hat{z}$. (4)
Here $\hat{\bm{k}}=\bm{k}/|k|$ and $F_{\omega_{r},k}^{a,p}$ are calculated microscopically using the Keldysh Green’s function formalism within the framework of the Usadel equations (See Sec. III.2 and Sec.III.4).
It is important to note that in the absence of spin–momentum locking, the superconducting order parameter does not couple linearly to magnons[55]. In S/FI hybrids based on conventional superconductors without SOC, only a linear-response triplet correction to the superconducting correlations arises due to magnon excitations. As a result, magnons can be dressed by triplet pairs forming magnon-cooparons [55], but composite excitations of magnons and superconducting collective modes do not occur because of the absence of a linear response singlet correction.
In the present case, the linear coupling arises directly from the fact that, in a TS, the superconducting correlations consist of mixed-spin components with equal singlet and triplet amplitudes [125]. A second key observation is that $F_{\omega_{r},k}^{a}=0$, as shown by microscopic calculations presented in App. B. This result implies that magnons do not excite the amplitude (Higgs) mode in a topological superconductor characterized by full spin–momentum locking and residing in its helical ground state.
On the other hand, the superconducting phase NG mode generates ac current response in the superconductor. The magnon-induced ac effective exchange field $\delta\bm{h}$ also produces ac electric current in the TS via the inverse magnetoelectric effect [126, 127, 128, 129, 130, 117, 131, 132, 133, 134, 135, 136]. The total ac current takes the form
$\bm{J}=(\bm{J}_{\omega_{r},k}e^{i\left(\omega_{r}t+\bm{k}\cdot\bm{r}\right)}+\bm{J}_{\omega_{r},\bm{k}}^{*}e^{-i\left(\omega_{r}t+\bm{k}\cdot\bm{r}\right)})e^{-\kappa t}.$ (5)
Up to the linear order with respect to the OP perturbations and $\delta\bm{h}$ it can be expanded as:
$\bm{J}_{\omega_{r},\bm{k}}=\bm{J}_{p}\delta\Delta^{p}+\bm{J}_{\delta h}\delta\bm{h}_{\omega,\bm{k}},$ (6)
[…]
$\bm{s}=-\frac{1}{2ev_{f}}(\hat{z}\times\bm{J}).$ (7)
The dynamics of spin wave in FI is described by the LLG equation:
In this equation $\bm{H_{eff}}=Km_{x}\hat{x}$, where $K$ is the uniaxial anisotropy constant, $D_{m}$ is the magnon stiffness and $\alpha$ is the Gilbert damping parameter. The last term represents the spin torque, which accounts for the back-action of the TS on the FI via the interfacial exchange interaction between the FI magnetization and the spin polarization $\bm{s}$ in TS. The factor $1/d_{f}$, where $d_{f}$ is the thickness of the FI layer, arises from averaging the interface exchange term over the FI thickness. Thus, in the TS, the NG superconducting mode generates an ac spin polarization, which in turn can excite magnons in the FI via the torque term in the LLG equation.
The nonzero response of the OP on the magnon, described by Eq. (4) together with the nonzero response of the FI magnetization on the NG mode described by the LLG equation (8) and Eq. (7) results in the appearance of composite excitations of magnons and the NG mode in FI/TS heterostructures. The hybrid excitations can be investigated in the basis $\hat{\Psi}=(\delta\Delta_{p},\delta h)^{T}$, where $\delta h$ is directly proportional to $\delta m_{y}$ via $\delta m_{y}=-2\gamma d_{s}\delta h/(J_{ex})M_{s}$. The spectrum of the hybridized excitations can be found from the combination of the linearized self-consistency equation for the OP and the linearized LLG equation, which can be written as follows (See Sec. IV.3 for the derivation and definition of $M^{p}$):
$\begin{pmatrix}M^{p}&M^{p}F^{p}\\ c_{s}(\bm{J}_{p}\cdot\hat{x})&M^{h}\end{pmatrix}\hat{\Psi}_{p}=0,$ (9)
The upper line of Eq. (9) is the linearized self-consistency equation for the superconducting gap and the bottom line comes from the LLG equation with $c_{s}=J_{ex}h_{0}/(ev_{f}d_{f})$ and $M^{h}=-(\omega_{b}+i\alpha[\omega_{r}+i\kappa])+c_{s}(\bm{J}_{\delta h}\cdot\hat{x})+[\omega_{r}+i\kappa]^{2}/(\omega_{b}+i\alpha\omega)$.
In the absence of the interface exchange coupling $J_{ex}=0$, that is $\delta h=0$, equation $M^{h}=0$ determines the bare magnon modes $\omega_{b}=\gamma K+D_{m}k^{2}$ and $M^{p}=0$ provides the phase NG mode of the charged 2D superconductor [72, 96, 98] with dispersion relation
$\hbar\omega_{NG}=\pi\sqrt{2\Delta_{0}k_{B}T_{c}\left(\frac{\hbar^{2}\omega_{p}^{2}k\xi}{4\pi^{2}\Delta_{0}k_{B}T_{c}}+(k\xi)^{2}\right)},$ (10)
where $\omega_{p}=\sqrt{4\pi^{2}e^{2}\nu\Delta_{0}D/\hbar\xi}=\sqrt{2\pi^{2}\sigma d_{s}\Delta_{0}/(\hbar\xi)}$ is the superconducting plasma frequency with $D$ being the diffusion constant of the TS. The spectrum of the NG mode is schematically shown in Fig. 1 (d).
Table 1: Problem description
MF potential Equation Observable
$\delta\Delta$ Gap equation Superconducting gap $\Delta$
$\delta\phi$ Poisson equation Electron density $\rho$
$\delta\bm{m}$ LLG equation Spin density $\bm{s}$
At $J_{ex}\neq 0$ the NG mode and the magnon are coupled due to nonzero $c_{s}$ and $F^{p}$ in Eq. (9). The coupled modes are represented in Fig. 1 (e). Due the spin-momentum locking which dictates the symmetry of the linear response in the form $(\hat{\bm{k}}\times\delta\bm{h}_{\omega,\bm{k}})\cdot\hat{z}$ the coupling is anisotropic. Its magnitude is maximal when the magnon propagates along the equilibrium magnetization $\hat{\bm{k}}=\hat{x}$, leading to the anticrossing between the NG and the magnon modes, and is zero if $\hat{\bm{k}}=\hat{y}$. Thus, the eigenmodes of the TS/FI bilayer become composite excitations magnon - Nambu-Goldstone. The analytical expressions for the anticrossing strength are derived in Sec. IV.3. However, even without doing microscopic calculations from Eq. (9) it is seen that the strength of interaction between the magnon and the NG mode, which can be quantified by $\delta\omega_{a}=\omega_{up}(k_{i})-\omega_{dn}(k_{i})$ (See Fig. 1 e), scales linearly with the interface exchange coupling $\delta\omega_{a}\propto\sqrt{M^{p}F^{p}c_{s}(\bm{J}_{p}\cdot\hat{x})}\propto h_{0}$ since $c_{s}\propto h_{0}^{2}$ and the other involved quantities do not contain $h_{0}$. The dependence of $\delta\omega_{a}$ on temperature is illustrated schematically in Fig. 1(f). It tends to zero at $T\to T_{c}$ indicating that the effect is of purely superconducting nature and strongly differs from the nonsuperconducting magnon-plasmon coupling of the magneto-electric origin, which was reported in [105, 106].